OpenStudy (anonymous):

i got a few questions to ask Evaluate 8 − 12 / 2 + 3 Evaluate 42 / (5 − 3) Evaluate (5 + 2)2 − 40 + 3 Evaluate 48 + 24 / 8 - 22 Help me with the 4 questions please

OpenStudy (anonymous):

Help me someone please :(

OpenStudy (anonymous):

1. -4/5 2.42/2=21/1=21 3.-23 4.72/-14=36/-7

OpenStudy (radar):

8-12/2+3=8-6+3=5 42/(5-3)=42/2=21 (5+2)2-40+3=14-40+3=-23 48+3-22=29

OpenStudy (radar):

5 21 -23 29